# Not flexible enough? Impacts of electric carsharing on a power sector with variable renewables11footnotemark: 1

### v1

@@ D.3 Power dispatch
Figure 16: Residual load curve, BEV charging and discharging for a subset of summer hours (upper panel) and winter hours (lower panel) [“Shared + other BEVs”, smart charging, high uptake]
Figure 17: Residual load curve, BEV charging and discharging for a subset of summer hours (upper panel) and winter hours (lower panel) [“Shared only”, bidirectional charging, high uptake]
D.4 Storage capacity and dispatch
@@
D.5 Industrial hydrogen demand
@@
Figure 28: Storage energy capacity in reference settings comparing without and with industrial demand for green hydrogen [“Shared + other BEVs”]
Figure 29: Storage energy capacity in carsharing scenarios comparing without and with industrial demand for green hydrogen [“Shared + other BEVs”]

### v2

@@ D.4 Power dispatch
Figure 18: Residual load curve, BEV charging for a subset of summer hours [“Shared + other BEVs”, smart charging, high uptake]
Figure 19: Residual load curve, BEV charging and discharging for a subset of summer hours [“Shared only”, bidirectional charging, high uptake]
D.5 Storage capacity and dispatch
@@
D.6 Sensitivity analysis: additional green hydrogen production
@@
D.7 Sensitivity analysis: weather years
Figure 30: System costs in reference settings for different weather years [“Shared + other BEVs”, high uptake]
Figure 31: Costs per substituted car in scenario settings for different weather years [“Shared + other BEVs”, high uptake]
Figure 32: Power generation capacity mixes in reference settings for different weather years [“Shared + other BEVs”, high uptake]
Figure 33: Power generation capacity mixes in scenario settings for different weather years [“Shared + other BEVs”, high uptake]
Figure 34: Yearly electricity generation in reference settings for different weather years [“Shared + other BEVs”, high uptake]
Figure 35: Yearly electricity generation in scenario settings for different weather years [“Shared + other BEVs”, high uptake]
Figure 36: Long-duration storage energy capacity in reference settings for different weather years [“Shared + other BEVs”, high uptake]
Figure 37: Long-duration storage energy capacity in scenario settings for different weather years [“Shared + other BEVs”, high uptake]
D.8 Sensitivity analysis: technology costs
Figure 38: System costs in reference settings for different technology cost assumptions [“Shared + other BEVs”, bidirectional charging, high uptake]
Figure 39: Costs per substituted car in scenario settings for different technology cost assumptions [“Shared + other BEVs”, bidirectional charging, high uptake]
Figure 40: Power generation capacity mixes in reference settings for different technology cost assumptions [“Shared + other BEVs”, bidirectional charging, high uptake]
Figure 41: Power generation capacity mixes in scenario settings for different technology cost assumptions [“Shared + other BEVs”, bidirectional charging, high uptake]
Figure 42: Yearly electricity generation in reference settings for different technology cost assumptions [“Shared + other BEVs”, bidirectional charging, high uptake]
Figure 43: Yearly electricity generation in scenario settings for different technology cost assumptions [“Shared + other BEVs”, bidirectional charging, high uptake]
D.9 Sensitivity analysis: driving electricity consumption
Smart charging Bidirectional charging
Reference (in billion euros) 46.0 44.9
Difference (in million euros) 306.6 676.4
Same trips Difference (in euros per substituted car) 49.2 108.4
Reference (in billion euros) 46.0 44.9
Difference (in million euros) 160.6 518.4
Shorter trips Difference (in euros per substituted car) 25.8 83.1
Table 11: Power sector costs for reference and scenarios [“Shared + other BEVs”, high uptake]. The “shorter trips” setting refers to a sensitivity analysis where trips undertaken with shared cars correspond to a 10% smaller driving electricity consumption compared to the main setting where mobility needs are assumed constant.
Figure 44: Generation capacity [“Shared + other BEVs”, high uptake]
Figure 45: Yearly electricity generation [“Shared + other BEVs”, high uptake]